OpenStudy (anonymous):

Determine the pH of a 0.260M solution of oxalic acid.

OpenStudy (abb0t):

pH = -log[\(H^+\)]

OpenStudy (anonymous):

So Ka isn't necessary?

OpenStudy (abb0t):

Yes, set up the expression of course for your acid. Remember its \(K_a =\large \frac{[product]}{[reactant]}\) can you set up your equation?

OpenStudy (anonymous):

So, like this?: \[H_{2}C _{2}O _{4}+H _{2}O \leftarrow \rightarrow HC _{2}O ^{-} _{4}+H _{3}O\] \[K _{a}=[HC _{2}O ^{-} _{4}][H _{3}O]/[H_{2}C _{2}O _{4}][H _{2}O]\]

OpenStudy (anonymous):

But then how do we utilize that? Can you walk me through please?

OpenStudy (abb0t):

For any acid, you generally have: \(HA \rightarrow H^+ + A^-\) so you'd have: \(\large \frac{[H^+][A^-]}{[HA]}\) DOES THAT make sens?

OpenStudy (anonymous):

No. Is what I have incorrecT?

OpenStudy (abb0t):

\(\LARGE\frac{[H^+][HC_2O_4^-]}{H_2C_2O_4}\)

OpenStudy (anonymous):

Hmm, okay. But the question says oxalic acid SOLUTION, so would the reactants not just be oxalic acid and water?

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Hi?

OpenStudy (aaronq):

you don't include water in equilibrium expressions

OpenStudy (aaronq):

you have to make an ICE table to figure the extend of the dissociation of the acid. When you you do that you'll find the value for H3O+, then just plug that into the formula abb0t wrote earlier

OpenStudy (anonymous):

Could you assist me with that please? I'm really struggling with what values to put in the ICE table

OpenStudy (aaronq):

yeah no problem, what do you have so far?

OpenStudy (anonymous):